OpenStudy (anonymous):

Which statement is true about exponential and logarithmic functions? Select one: a. The domain of the exponential function is the domain of the logarithmic function. b. The y values of the exponential function are the y values of the logarithmic function. c. The x values of the logarithmic function are the y values of the exponential function. d. All of the above.

OpenStudy (anonymous):

@BSwan

OpenStudy (anonymous):

let's say your y = e^x to isolate x, you use natural log ln y = x so compare y = e^x (exponential) and ln y = x (logarithmic)

OpenStudy (anonymous):

so which one would it be

OpenStudy (anonymous):

It's obvious if you "read" the last line I wrote

OpenStudy (anonymous):

I got B is that right.. I just don't want to meess up

OpenStudy (anonymous):

No. Their graphs look very different. It's C

OpenStudy (anonymous):

ok thanks can you help on somemore please